The perimeter of a rectangle is 48 meters. The length is 18 meters less than six times width. Find the dimensions of the rectangle.
Answer by macston(5194)

.
W=width; L=length=6W-18
.
P=2(L+W)
48m=2(L+W)
24m=L+W
24m=(6W-18)+W
24m=7W-18
42m=7W
6=W
ANSWER 1: The width is 6 meters.
.
L=6W-18=6(6)-18=36-18=18
ANSWER 2: The length is 18 meters.
.
CHECK:
.
P=2(L+W)
48m=2(18m+6m)
48m=2(24m)
48m=48m
.